OpenStudy (azionne):

7(2x + 6) = 1 will medal

pooja195 (pooja195):

\[\huge~\rm~(7)(2x)+(7)(6)=1\]

OpenStudy (azionne):

this would be 14x=42 right

pooja195 (pooja195):

14x+42=1

OpenStudy (anonymous):

u must start this @Azionne

pooja195 (pooja195):

Then subtract 42 from both sides

OpenStudy (anonymous):

good start :) just a silly mistake 14x+42=1

pooja195 (pooja195):

14x+42-42=1-42 14x=1-42

pooja195 (pooja195):

1-42=?

OpenStudy (azionne):

14x=-41

OpenStudy (azionne):

or 14x=41

OpenStudy (anonymous):

perfect :)

pooja195 (pooja195):

Good next just divide both sides by 14 14x/14=-41/14

OpenStudy (anonymous):

its negatvi Smaller - Greater = - ans

OpenStudy (anonymous):

negative**

OpenStudy (azionne):

-2.928571428571429

pooja195 (pooja195):

Keep it in fraction form x=-41/14

OpenStudy (azionne):

oh ok thx

pooja195 (pooja195):

np

OpenStudy (mathmale):

Actually, it makes sense, and is much more accurate, to leave your answer in fraction form, as Pooja has pointed out.

OpenStudy (anonymous):

u learned it :)

OpenStudy (mathmale):

Usually long division leads to round-off errors.

OpenStudy (azionne):

yes i did still a little rusty but will get soon

pooja195 (pooja195):

That's understandable jut takes practice =)

OpenStudy (anonymous):

usually keep ur till 4 decimal places

OpenStudy (anonymous):

up**